Use a comma to separate answers as needed, unless it is not a real number
Solve: The square root of 5x + 29 = x + 3
Thanks so much!!

Answer by nyc_function(2741)

sqrt{5x + 29} = x + 3
Square both sides to begin. By squaring both sides, you will then be able to remove the radical from the left side.
[sqrt{5x + 29}]^2 = (x + 3)^2
5x + 29 = x^2 + 6x + 9

0 = x^2 + 6x - 5x + 9 - 29
0 = x^2 + x - 20
0 = (x - 4)(x + 5)
Set each factor to zero and solve for x.
x - 4 = 0
x = 4
OR
x + 5 = 0
x = -5
To make sure that x = 4 and x = -5 are correct, plug those two values for x into the original radical equation given. The correct x value will produce the same answer on both sides of the equation after simplifying. I'll let you do that, however.